# Carrier Aggregation Enabled MIMO-OFDM Integrated Sensing and Communication

### v1

@@ III-B2 Fusion algorithm of Doppler feature vectors
Without considering the noise, observing (30), the different parameters between $\mathbf{V}_{S}^{1}$ and $\mathbf{V}_{S}^{2}$ are carrier frequency and total symbol duration. (30) is rewritten as
$\mathbf{V}_{S}^{b}=\left[e^{j2\pi m\frac{2v_{0}}{c}\left(f_{\mathrm{C}}^{b}T^{b}\right)}\right]|_{m=0,1,\cdots,M_{b}-1},$ (36)
where $f_{\mathrm{C}}^{b}T^{b}=f_{\mathrm{C}}^{b}\left(\frac{1}{\Delta f^{b}}+T_{\text{s}}^{b}\right)$. If $f_{\mathrm{C}}^{1}T^{1}=f_{\mathrm{C}}^{2}T^{2}$, $\mathbf{V}_{S}^{1}$ and $\mathbf{V}_{S}^{2}$ can be coherently accumulated to obtain a SNR gain.
@@
The fusion process of Doppler feature vectors is revealed in the following three cases.
Case 1 $(M_{1}>M_{2})$: Initialize a zero-element vector $\mathbf{E}_{S}^{1}\in\mathbb{C}^{1\times M_{1}}$ and complement $\mathbf{V}_{S}^{2}$ with zero to get the matrix $\hat{\mathbf{V}}_{S}^{2}\in\mathbb{C}^{1\times M_{1}}=[\mathbf{V}_{S}^{2},0_{M_{2}},\cdots,0_{M_{1}-1}]$. Then, the processed vector $\mathbf{E}_{S}^{1}=(\mathbf{V}_{S}^{1}+\hat{\mathbf{V}}_{S}^{2})/2$ is obtained.
Case 2 $(M_{1}<M_{2})$: Initialize a zero-element vector $\mathbf{E}_{S}^{2}\in\mathbb{C}^{1\times M_{2}}$ and complement $\mathbf{V}_{S}^{1}$ with zero to get the matrix $\hat{\mathbf{V}}_{S}^{1}\in\mathbb{C}^{1\times M_{2}}=[\mathbf{V}_{S}^{1},0_{M_{1}},\cdots,0_{M_{2}-1}]$. Then, the processed vector $\mathbf{E}_{S}^{2}=(\hat{\mathbf{V}}_{S}^{1}+\mathbf{V}_{S}^{2})/2$ is obtained.
Case 3 $(M_{1}=M_{2}=M)$: Initialize a zero-element vector $\mathbf{E}_{S}^{3}\in\mathbb{C}^{1\times M}$ and obtain a vector $\mathbf{E}_{S}^{3}=(\mathbf{V}_{S}^{1}+\mathbf{V}_{S}^{2})/2$.
Taking the $\mathbf{E}_{S}^{1}$ in Case 1 as an example, an improved discrete Fourier transform (DFT) method is applied to achieve high-accuracy sensing. The specific procedures are as follows, while the Case 2 and Case 3 have the similar procedures.
@@
Step 2: Generate a velocity searching vector $\vec{\mathbf{B}}$ by gridding the searching interval with a size of grid being $\Delta V=\frac{V_{\text{max}}-V_{\text{min}}}{G}$, where $G$ is the number of grids and $\vec{\mathbf{B}}\in\mathbb{C}^{1\times G}$ is expressed as
$\vec{\mathbf{B}}=\left[V_{1},V_{2},\cdots,V_{g},\cdots,V_{G}\right],$ (40)
@@
Step 3: According to the size of $\mathbf{E}_{S}^{1}$ and the expression form of Doppler feature vectors in (30), $\vec{\mathbf{B}}$ is transformed to a searching matrix $\mathbf{B}_{1}\in\mathbb{C}^{M_{1}\times G}$, which is expressed as
@@
Step 4: Obtain and search the Doppler profile $\vec{\mathbf{E}}_{1}=\mathbf{E}_{S}^{1}\mathbf{B}_{1}$. The peak index of $\vec{\mathbf{E}}_{1}$ is denoted by $\hat{g}_{1}$.
Step 5: The velocity estimation of target is $\hat{v}_{0}^{1}=\vec{\mathbf{B}}(\hat{g}_{1})$.
The fusion algorithm of Doppler feature vectors is shown in Algorithm 2, which is explained intuitively in Fig. 4.
@@
Doppler feature vectors $\mathbf{V}_{S}^{1}$ and $\mathbf{V}_{S}^{2}$;
@@
Output: The velocity estimations of target $\hat{v}_{0}^{1}$, $\hat{v}_{0}^{2}$, and $\hat{v}_{0}^{3}$.
@@
2:     Assume a vector $\mathbf{E}_{S}^{1}\in\mathbb{C}^{1\times M_{1}}=\mathbf{V}_{S}^{1}$,
@@
3:     for $\mathbf{E}_{S}^{1}$ $\omega_{1}$ in $\mathcal{M}_{1}$ do
4:        The element value of $\mathbf{V}_{S}^{2}(\omega_{1})$ is added to $\mathbf{E}_{S}^{1}(\omega_{1})$;
5:        $\mathbf{E}_{S}^{1}(\omega_{1})=\mathbf{E}_{S}^{1}(\omega_{1})/2$;
@@
8:     Assume a vector $\mathbf{E}_{S}^{2}\in\mathbb{C}^{1\times M_{2}}=\mathbf{V}_{S}^{2}$,
@@
9:     for $\mathbf{E}_{S}^{2}$ $\omega_{2}$ in $\mathcal{M}_{2}$ do
10:        The element value of $\mathbf{V}_{S}^{1}(\omega_{2})$ is added to $\mathbf{E}_{S}^{2}(\omega_{2})$;
11:        $\mathbf{E}_{S}^{2}(\omega_{2})=\mathbf{E}_{S}^{2}(\omega_{2})/2$;
@@
14:     Initialize a zero-elements vector $\mathbf{E}_{S}^{3}\in\mathbb{C}^{1\times M}$,
@@
15:     for $\mathbf{E}_{S}^{3}$ $\omega_{3}$ in $\mathcal{M}_{3}$ do
16:        $\mathbf{E}_{S}^{3}(\omega_{3})=\left[\mathbf{V}_{S}^{1}(\omega_{3})+\mathbf{V}_{S}^{2}(\omega_{3})\right]/2$;
@@
19: The results of $\mathbf{E}_{S}^{1}$, $\mathbf{E}_{S}^{2}$, and $\mathbf{E}_{S}^{3}$ undergoing improved DFT are
@@
20: Substitute $\hat{g}_{1}$, $\hat{g}_{2}$, and $\hat{g}_{3}$ into (40) to obtain the velocity
estimations of target $\hat{v}_{0}^{1}$, $\hat{v}_{0}^{2}$, and $\hat{v}_{0}^{3}$.
@@ IV-A1 Communication MI of MIMO-OFDM ISAC system
According to the signal model in (42), the communication MI of the MIMO-OFDM ISAC system is obtained in (43), with $N_{x}$, $N_{c}$, and $N_{r}$ denoting the number of OFDM symbols, subcarriers, and received antennas, respectively [27].
@@ IV-A2 Communication MI of the proposed CA-enabled MIMO-OFDM ISAC system
$I\left(\mathbf{X}_{\mathrm{C}}^{1},\mathbf{X}_{\mathrm{C}}^{2};\mathbf{Y}_{u}^{\mathrm{C}}\mid\mathbf{H}_{\mathrm{C}}^{1},\mathbf{H}_{\mathrm{C}}^{1}\right).$ (44)
@@
For ease of derivation, we consider the assumptions of $N_{1}=N_{2}=N$ and $M_{1}=M_{2}=M$. Since $\mathbf{X}_{\mathrm{C}}^{b}$ is a block diagonal array, the probability density function (PDF) of $\mathbf{Y}_{u}^{\mathrm{C}}$ under known CSI is the product of the PDFs on the $N$ subcarriers, denoted by (47), where $\mathbf{y}_{u}^{\mathrm{C},j}$ is the $j$-th row or column vector of $\mathbf{Y}_{u}^{\mathrm{C}}$, and $\mathbf{I}_{N_{\text{U}}}$ represents a $N_{\text{U}}\times N_{\text{U}}$ unit matrix.
@@ IV-B Sensing Performance Analysis
The CRLB is generally used to determine the lower bound on the variance of unbiased estimator and characterize the sensing ability of a signal [28]. Therefore, the CRLB of the CA-enabled MIMO-OFDM ISAC system is derived in this section.
@@ IV-B1 CRLB of MIMO-OFDM system
The CRLB for OFDM ISAC system is derived in [7], and the CRLB for MIMO-OFDM ISAC system is revealed in Theorem 3. For ease of derivation, $\gamma=\frac{2v_{0}}{c}$ is used as an estimation of velocity to avoid the impact of different carrier frequency on theoretical derivation [7]. Meanwhile, $\sin{\theta_{\text{Rx}}}$ is used as an estimation of AoA to avoid the presence of unknown parameters after multiple derivations.
@@
For a MIMO-OFDM ISAC system (e.g., the $b$-th CC), the CRLBs of $\sin{\theta_{\text{Rx}}}$, $\tau_{0}$, and $\gamma$ are expressed as (52), (53), and (54), respectively.
$CRLB(\sin{\theta_{\text{Rx}}})=\frac{\Gamma_{\text{mimo}}N_{\text{U}}\left[M_{b}N_{b}\Theta_{b}^{n}\Theta_{b}^{m}-\left(\Theta_{b}^{m,n}\right)^{2}\right]}{M_{b}N_{b}},$ (52)
$CRLB(\tau_{0})=\frac{\Gamma_{\text{mimo}}N_{b}\left[M_{b}N_{\text{U}}\Theta_{b}^{p}\Theta_{b}^{m}-\left(\Theta_{b}^{p,m}\right)^{2}\right]}{M_{b}N_{\text{U}}},$ (53)
$CRLB(\gamma)=\frac{\Gamma_{\text{mimo}}M_{b}\left[N_{b}N_{\text{U}}\Theta_{b}^{p}\Theta_{b}^{n}-\left(\Theta_{b}^{p,n}\right)^{2}\right]}{N_{b}N_{\text{U}}},$ (54)
@@
$\Gamma_{\text{mimo}}=\frac{1}{{\varpi\cdot\left(2\pi\right)^{2}}}\frac{1}{M_{b}N_{b}N_{\text{U}}\Theta_{b}^{p}\Theta_{b}^{n}\Theta_{b}^{m}+2\Theta_{b}^{p,n}\Theta_{b}^{p,m}\Theta_{b}^{m,n}-N_{b}\left(\Theta_{b}^{p,m}\right)^{2}\Theta_{b}^{n}-M_{b}\left(\Theta_{b}^{p,n}\right)^{2}\Theta_{b}^{m}-N_{\text{U}}\left(\Theta_{b}^{m,n}\right)^{2}\Theta_{b}^{p}},$ (55)

### v2

@@ III-B2 Fusion algorithm of Doppler feature vectors
Taking the $\mathbf{e}_{i^{\prime},S}^{1}$ and $\mathbf{e}_{i^{\prime},S}^{2}$ as an example. Without considering the interference, observing (25), the different parameters between $\mathbf{e}_{i^{\prime},S}^{1}$ and $\mathbf{e}_{i^{\prime},S}^{2}$ are carrier frequency and total symbol duration. (25) is rewritten as
$\mathbf{e}_{i^{\prime},S}^{b}=\left[e^{j2\pi m\frac{2v_{i^{\prime},0}}{c}\left(f_{\mathrm{C}}^{b}T^{b}\right)}\right]|_{m=0,1,\cdots,M_{b}-1},$ (31)
where $f_{\mathrm{C}}^{b}T^{b}=f_{\mathrm{C}}^{b}\left(\frac{1}{\Delta f^{b}}+T_{\text{s}}^{b}\right)$. If $f_{\mathrm{C}}^{1}T^{1}=f_{\mathrm{C}}^{2}T^{2}$, $\mathbf{e}_{i^{\prime},S}^{1}$ and $\mathbf{e}_{i^{\prime},S}^{2}$ can be coherently accumulated to obtain a SNR gain.
@@
Similar to the fusion of delay feature vectors, the $\tilde{\mathbf{v}}_{i^{\prime},S}^{1}=\frac{\sigma_{S,2}^{2}}{\sigma_{S,1}^{2}+\sigma_{S,2}^{2}}\mathbf{e}_{i^{\prime},S}^{1}$ and $\tilde{\mathbf{v}}_{i^{\prime},S}^{2}=\frac{\sigma_{S,1}^{2}}{\sigma_{S,1}^{2}+\sigma_{S,2}^{2}}\mathbf{e}_{i^{\prime},S}^{2}$ are obtained. The fusion process of Doppler feature vectors is revealed in the following three cases.
Case 1 $(M_{1}>M_{2})$: Initialize a zero-element vector $\mathbf{f}_{i^{\prime},S}^{1}\in\mathbb{C}^{1\times M_{1}}$ and complement $\tilde{\mathbf{v}}_{i^{\prime},S}^{2}$ with zero to get the matrix $\hat{\mathbf{v}}_{i^{\prime},S}^{2}\in\mathbb{C}^{1\times M_{1}}=[\tilde{\mathbf{v}}_{i^{\prime},S}^{2},0_{M_{2}},\cdots,0_{M_{1}-1}]$. Then, the processed vector $\mathbf{f}_{i^{\prime},S}^{1}=(\tilde{\mathbf{v}}_{i^{\prime},S}^{1}+\hat{\mathbf{v}}_{i^{\prime},S}^{2})/2$ is obtained.
Case 2 $(M_{1}<M_{2})$: Initialize a zero-element vector $\mathbf{f}_{i^{\prime},S}^{2}\in\mathbb{C}^{1\times M_{2}}$ and complement $\tilde{\mathbf{v}}_{i^{\prime},S}^{1}$ with zero to get the matrix $\hat{\mathbf{v}}_{i^{\prime},S}^{1}\in\mathbb{C}^{1\times M_{2}}=[\tilde{\mathbf{v}}_{i^{\prime},S}^{1},0_{M_{1}},\cdots,0_{M_{2}-1}]$. Then, the processed vector $\mathbf{f}_{i^{\prime},S}^{2}=(\hat{\mathbf{v}}_{i^{\prime},S}^{1}+\tilde{\mathbf{v}}_{i^{\prime},S}^{2})/2$ is obtained.
Case 3 $(M_{1}=M_{2}=M)$: Initialize a zero-element vector $\mathbf{f}_{i^{\prime},S}^{3}\in\mathbb{C}^{1\times M}$ and obtain a vector $\mathbf{f}_{i^{\prime},S}^{3}=(\tilde{\mathbf{v}}_{i^{\prime},S}^{1}+\tilde{\mathbf{v}}_{i^{\prime},S}^{2})/2$.
Taking the $\mathbf{E}_{i^{\prime},S}^{1}$ in Case 1 as an example, an improved discrete Fourier transform (DFT) method is applied to achieve high-accuracy sensing. The specific procedures are as follows, while the Case 2 and Case 3 have the similar procedures.
@@
Step 2: Generate a velocity searching vector $\mathbf{b}$ by gridding the searching interval with a size of grid being $\Delta V=\frac{V_{\text{max}}-V_{\text{min}}}{G}$, where $G$ is the number of grids and $\mathbf{b}\in\mathbb{C}^{1\times G}$ is expressed as
$\mathbf{b}=\left[V_{1},V_{2},\cdots,V_{g},\cdots,V_{G}\right],$ (35)
@@
Step 3: According to the size of $\mathbf{f}_{i^{\prime},S}^{1}$ and the expression form of Doppler feature vectors in (25), $\mathbf{b}$ is transformed to a searching matrix $\mathbf{B}_{1}\in\mathbb{C}^{M_{1}\times G}$, which is expressed as
@@
Step 4: Obtain and search the Doppler profile $\tilde{\mathbf{f}}_{i^{\prime},1}=\mathbf{f}_{i^{\prime},S}^{1}\mathbf{B}_{1}$. The peak index of $\tilde{\mathbf{f}}_{i^{\prime},1}$ is denoted by $\hat{g}_{i^{\prime}}^{1}$.
Step 5: The velocity estimation of the $i^{\prime}$-th target is $\hat{v}_{i^{\prime},0}^{1}=\mathbf{b}(\hat{g}_{i^{\prime}}^{1})$.
The fusion algorithm of Doppler feature vectors for the $i^{\prime}$-th target is shown in Algorithm 2, which is explained intuitively in Fig. 4. Finally, the estimated velocities of $I$ targets are obtained.
@@
Doppler feature vectors of the $i^{\prime}$-th target $\tilde{\mathbf{v}}_{i^{\prime},S}^{1}$ and $\tilde{\mathbf{v}}_{i^{\prime},S}^{2}$;
@@
Output: The velocity estimations of the $i^{\prime}$-th target $\hat{v}_{i^{\prime},0}^{1}$, $\hat{v}_{i^{\prime},0}^{2}$, and $\hat{v}_{i^{\prime},0}^{3}$.
@@
2:     Assume a vector $\mathbf{f}_{i^{\prime},S}^{1}\in\mathbb{C}^{1\times M_{1}}=\tilde{\mathbf{v}}_{i^{\prime},S}^{1}$,
@@
3:     for $\omega_{1}$ in $\mathcal{M}_{1}$ do
4:        The element value of $\tilde{\mathbf{v}}_{i^{\prime},S}^{2}(\omega_{1})$ is added to $\mathbf{f}_{i^{\prime},S}^{1}(\omega_{1})$;
5:        $\mathbf{f}_{i^{\prime},S}^{1}(\omega_{1})=\mathbf{f}_{i^{\prime},S}^{1}(\omega_{1})/2$;
@@
8:     Assume a vector $\mathbf{f}_{i^{\prime},S}^{2}\in\mathbb{C}^{1\times M_{2}}=\tilde{\mathbf{v}}_{i^{\prime},S}^{2}$,
@@
9:     for $\omega_{2}$ in $\mathcal{M}_{2}$ do
10:        The element value of $\tilde{\mathbf{v}}_{i^{\prime},S}^{1}(\omega_{2})$ is added to $\mathbf{f}_{i^{\prime},S}^{2}(\omega_{2})$;
11:        $\mathbf{f}_{i^{\prime},S}^{2}(\omega_{2})=\mathbf{f}_{i^{\prime},S}^{2}(\omega_{2})/2$;
@@
14:     Initialize a zero-elements vector $\mathbf{f}_{i^{\prime},S}^{3}\in\mathbb{C}^{1\times M}$,
@@
15:     for $\omega_{3}$ in $\mathcal{M}_{3}$ do
16:        $\mathbf{f}_{i^{\prime},S}^{3}(\omega_{3})=\left[\tilde{\mathbf{v}}_{i^{\prime},S}^{1}(\omega_{3})+\tilde{\mathbf{v}}_{i^{\prime},S}^{2}(\omega_{3})\right]/2$;
@@
19: The results of $\mathbf{f}_{i^{\prime},S}^{1}$, $\mathbf{f}_{i^{\prime},S}^{2}$, and $\mathbf{f}_{i^{\prime},S}^{3}$ undergoing improved DFT are
@@
20: Substitute $\hat{g}_{i^{\prime}}^{1}$, $\hat{g}_{i^{\prime}}^{2}$, and $\hat{g}_{i^{\prime}}^{3}$ into (35) to obtain the velocity
estimations of the $i^{\prime}$-th target $\hat{v}_{i^{\prime},0}^{1}$, $\hat{v}_{i^{\prime},0}^{2}$, and $\hat{v}_{i^{\prime},0}^{3}$.
@@ IV-A1 Communication MI of MIMO-OFDM ISAC system
According to the signal model in (37), the communication MI of the MIMO-OFDM ISAC system is obtained in (38), with $N_{x}$, $N_{c}$, and $N_{r}$ denoting the number of OFDM symbols, subcarriers, and receive antennas, respectively [26].
@@ IV-A2 Communication MI of the proposed CA-enabled MIMO-OFDM ISAC system
$I\left(\mathbf{X}_{\mathrm{C}}^{1},\mathbf{X}_{\mathrm{C}}^{2};\mathbf{Y}_{u}^{\mathrm{C}}\mid\mathbf{H}_{\mathrm{C}}^{1},\mathbf{H}_{\mathrm{C}}^{2}\right).$ (39)
@@
Without loss of generality, we consider the assumptions of $N_{1}=N_{2}=N$ and $M_{1}=M_{2}=M$. Since $\mathbf{X}_{\mathrm{C}}^{b}$ is a block diagonal array, the probability density function (PDF) of $\mathbf{Y}_{u}^{\mathrm{C}}$ under known CSI is the product of the PDFs on the $N$ subcarriers, denoted by (42), where $\mathbf{y}_{u}^{\mathrm{C},j}$ is the $j$-th row or column vector of $\mathbf{Y}_{u}^{\mathrm{C}}$, and $\mathbf{I}_{N_{\text{U}}}$ represents a $N_{\text{U}}\times N_{\text{U}}$ unit matrix.
@@
Upon observing (41), we find that the maximum communication channel capacity of the proposed ISAC system is affected by space-time-frequency resources and increases with the increase of the number of CC. the MI simulation analysis is given in Section V-A to verify the advantage of large channel capacity compared with MIMO-OFDM system.
@@ IV-B Sensing Performance Analysis
The CRLB is generally used to determine the lower bound on the variance of unbiased estimator and characterize the sensing ability of a signal [42]. Therefore, the CRLB of the proposed CA-enabled MIMO-OFDM ISAC system is derived in this section.
@@ IV-B1 CRLB of MIMO-OFDM system
The CRLB for the OFDM ISAC system has been derived in [8], and the CRLB for the MIMO-OFDM ISAC system is presented in Theorem 3. For ease of derivation, $\gamma_{i}=\frac{2v_{i,0}}{c}$ is used as an estimate of velocity to avoid the impact of varying carrier frequencies on theoretical derivation [8]. Meanwhile, $\sin{\theta_{i,\text{Rx}}}$ is used as an estimate of AoA to avoid the presence of unknown parameters after multiple derivations [43].
@@
For a MIMO-OFDM ISAC system (e.g., the $b$-th CC), the CRLBs of $\sin{\theta_{i,\text{Rx}}}$, $\tau_{i,0}$, and $\gamma_{i}$ are expressed as (47), (48), and (49), respectively.
$CRLB(\sin{\theta_{i,\text{Rx}}})=\frac{\Gamma_{\text{mimo}}N_{\text{R}}\left[M_{b}N_{b}\Theta_{b}^{n}\Theta_{b}^{m}-\left(\Theta_{b}^{m,n}\right)^{2}\right]}{M_{b}N_{b}},$ (47)
$CRLB(\tau_{i,0})=\frac{\Gamma_{\text{mimo}}N_{b}\left[M_{b}N_{\text{R}}\Theta_{b}^{p}\Theta_{b}^{m}-\left(\Theta_{b}^{p,m}\right)^{2}\right]}{M_{b}N_{\text{R}}},$ (48)
$CRLB(\gamma_{i})=\frac{\Gamma_{\text{mimo}}M_{b}\left[N_{b}N_{\text{R}}\Theta_{b}^{p}\Theta_{b}^{n}-\left(\Theta_{b}^{p,n}\right)^{2}\right]}{N_{b}N_{\text{R}}},$ (49)
@@
$\Gamma_{\text{mimo}}=\frac{1}{{\varpi\cdot\left(2\pi\right)^{2}}}\frac{1}{M_{b}N_{b}N_{\text{R}}\Theta_{b}^{p}\Theta_{b}^{n}\Theta_{b}^{m}+2\Theta_{b}^{p,n}\Theta_{b}^{p,m}\Theta_{b}^{m,n}-N_{b}\left(\Theta_{b}^{p,m}\right)^{2}\Theta_{b}^{n}-M_{b}\left(\Theta_{b}^{p,n}\right)^{2}\Theta_{b}^{m}-N_{\text{R}}\left(\Theta_{b}^{m,n}\right)^{2}\Theta_{b}^{p}},$ (50)